# reference.bib

### v2

@@ III PERFORMANCE EVALUATION OF 3D LANE DETECTION
This section will discuss the performance evaluation of monocular 3D lane detection models. Herein, we explain the evaluation metrics, different types of objective functions, analyze the computational complexity, and finally provide quantitative comparisons of various models. The nomenclature of the used variables is given in Table I.
First, we present the visualization results of 3D lane line detection. As some algorithms are not open-sourced, we have only performed visualization tests on the ApolloSim dataset using a few open-source algorithms. These algorithms have been trained on the ApolloSim dataset, and the visualization results are shown in Fig. 8, where the red lines represent the predicted lane lines, and the blue lines represent the ground truth lane lines.
Next, we will introduce the evaluation metrics, the loss function used for training the algorithms, and the quantitative test results of 3D lane line detection on public datasets.
@@ III-A Evaluation Metrics for 3D Lane Detection
Building only predictive monoculdar 3D lane detection models is not a wise and trustworthy decision for safe AD unless it is tested on unseen data. Most models evaluate their performance on a disjoint set of the same dataset used for training, i.e., the test data remains new to the training model. Deep learning models for monocular 3D lane detection are evaluated using some general metrics to assess the best results based on true values. For monocular 3D lane detection tasks, different types of evaluation metrics are available for these tasks, which we review next as follows:
@@ III-A7 Z Error
III-A8 CD
$\begin{cases}CD=\frac{1}{N}\sum_{i=1}^{N}||P_{gi}-\hat{P_{gj}}||_{2},\\ \hat{P_{pj}}=\min_{P_{p_{j}}\in L^{p}}||P_{pj}-P_{gi}||_{2}.\end{cases}$ (8)
@@ III-C Quantitative Analysis of Monocular 3D Lane Detection Models
This section elaborates on the quantitative empirical analysis of monocular 3D lane detection methods surveyed in this paper, which contribute to the automation of autonomous driving. For quantitative assessment, we utilize four evaluation metrics to examine the performance of each monocular 3D lane detection method on the ApolloSim dataset: AP, F-Score, x error and z error, and we report the results in Table IV. On the Openlane dataset, we evaluated the F-Score of each model, as shown in Table V. On the ONCE-3DLane dataset, we evaluate four metrics, namely: F-Score, Precision, Recall, and CD error, the results are reported om Table VI.
Additionally, computational efficiency is taken into account by reporting the Frames Per Second (FPS) achievable by each method during the inference process. The overall runtime of these models is reported in Table IV,V and VI. In some papers, the inference time of the algorithms and their corresponding hardware platforms are reported, which we directly utilize. However, in some other papers, the inference time of the algorithms is not shown, so we conduct our own experiments to test the inference time on our experimental platform. Our experimental platform’s CPU configuration includes an Intel(R) Core i9-12900K CPU processor running on the Ubuntu 20.04 operating system, while the GPU used in the experiments is an NVIDIA GeForce RTX 3080Ti GPU with 12 GB of graphics memory, In the above tables, we specified the hardware used for inference for each method.
@@ IV DATASETS
In vision tasks based on deep learning, another equally important component is the dataset. In this section, we will introduce the datasets currently available for monocular 3D lane line detection tasks. Some of these datasets are open-sourced and popular, frequently used by the community, while others are only described in papers and not open-sourced. Whether open-source or proprietary datasets, in order to provide a more intuitive overview of these datasets, we have compiled a detailed table showcasing all existing monocular 3D lane line detection datasets, as presented in Table III.
@@ IV-A Apollo 3D Lane Synthetic Dataset
TABLE II: A detailed description of some representative monocular 3D lane detection algorithms, including methodology, open-sourcing status, utilized datasets, and network architecture.
Ref. Year Methodology Dataset Code Network Architecture   Need BEV Trans.?
[garnett20193d] 2018 Combining novel network architecture, IPM, and anchor-based lane representation to achieve accurate 3D lane detection. 3D-Lanes, Synthetic-3D-Lanes N/A CNN                 ✔
[guo2020gen] 2020 Using a two-stage framework that decouples the learning of the image segmentation subnetwork and the geometry encoding subnetwork, reducing the amount of 3D lane labels required for training. Apollo Synthetic. link CNN                 ✔
[efrat2020semi] 2020 Introducing uncertainty estimation, utilizing a semi-local, Bird’s Eye View (BEV), tile representation to break down lanes into simple segments, and subsequently clustering them into complete lanes. Synthetic-3D-Lanes, 3D-lanes. N/A CNN                 ✔
[efrat20203d] 2020 An anchor-free method, which can detect complex lane topologies such as splits, merges, short lanes, and perpendicular lanes. Synthetic-3D-Lanes, Real-3D-lanes N/A CNN                 ✔
[liu2022learning] 2021 Presenting a two-stage framework, leveraging auxiliary 3D lane tasks and geometry constraints in the first stage, followed by accurate 3D lane prediction in the second stage using estimated pose. Apollo Synthetic. link CNN                 ✔
[chen2022persformer] 2022 This paper introduces a novel Transformer-based spatial feature transformation module and presents OpenLane, a large-scale real-world 3D lane dataset. Apollo Synthetic, OpenLane. link Transformer           ✔
[yan2022once] 2022 This paper introduces the ONCE-3DLanes dataset and proposes SALAD, an extrinsic-free, anchor-free method for regressing 3D lane coordinates in image view. ONCE-3DLanes. link CNN                 ✘
[jiang2022robust] 2022 This paper proposes Att-Gen-LaneNet, a two-stage network for robust 3D lane detection. It integrates the Efficient Channel Attention (ECA) and Convolutional Block Attention Module (CBAM) to improve segmentation and generalization. Apollo Synthetic. N/A CNN                 ✔
[li2022reconstruct] 2022 This paper proposes an approach for monocular 3D lane detection by leveraging the geometry structure and explicit supervision based on the structure prior. Apollo Synthetic. N/A CNN                 ✔
[li2022perspective] 2022 The method comprises three stages: generating multiscale front-view features, calculating bird’s eye view (BEV) features, and predicting 2D and 3D lanes. OpenLane. N/A CNN                 ✔
[bai2022curveformer] 2022 The method formulates 3D lane detection as a curve propagation problem, enhancing performance by leveraging curve queries, a curve cross-attention module, and a context sampling module to capture relevant image features. Apollo Synthetic, OpenLane. N/A Transformer           ✘
[ai2022ws] 2022 This paper introduces a method training 3D lanes with only 2D lane labels, leveraging assumptions of constant lane width and equal height on adjacent lanes for indirect supervision of 3D lane heights. Apollo Synthetic, ONCE-3DLanes. N/A CNN                 ✔
[wang2023bevlanedet] 2022 This paper introduces a Virtual Camera for spatial relationship consistency, a Key-Points Representation module for diverse 3D lane structures, and a Spatial Transformation Pyramid for front-view to bird’s eye view feature transformation. Apollo Synthetic, OpenLane. link CNN                 ✔
[wang2022spatio] 2022 Utilizing multi-frame input and removing the dependency on camera pose, the model integrates a multi-frame pre-alignment layer, spatio-temporal attention module, and 3DLane IOULoss. OpenLane, ONCE-3DLanes. N/A CNN                 ✔
[huang2023anchor3dlane] 2023 This paper introduces Anchor3DLane, a BEV-free method predicting 3D lanes directly from front-viewed (FV) representations using defined 3D lane anchors. Apollo Synthetic, OpenLane, ONCE-3DLanes. link CNN                 ✘
[kim2023d] 2023 This paper introduces a method that employs depth-aware voxel mapping, extends contextual features to 3D space, and predicts confidence intervals using Laplace loss. LLAMAS-3DLD, Pandaset-3DLD, Apollo Synthetic. N/A CNN                 ✔
[chen2023efficient] 2023 This paper proposes an efficient transformer for 3D lane detection using a decomposed cross-attention mechanism to learn lane and bird’s-eye-view (BEV) representations concurrently. OpenLane, ONCE-3DLanes. N/A Transformer           ✘
[li2023grouplane] 2023 Similar to Paper [qin2020ultra], this is the first work on 3D lane detection utilizing fully convolutional heads based on row-wise classification. OpenLane, ONCE-3DLanes, OpenLane-Huawei. N/A CNN                 ✔
[luo2023latr] 2023 This paper introduces LATR, an end-to-end 3D lane detection model utilizing 3D-aware front-view features and cross-attention mechanism with lane-aware query generator. OpenLane, Apollo Synthetic, ONCE-3DLanes. link Transformer           ✘
[han2023decoupling] 2023 This paper introduces a novel lane detection method, dividing it into curve modeling and ground height regression, within a unified framework for both 2D and 3D lane detection. OpenLane, ONCE-3DLanes, TuSimple, CULane. N/A CNN                 ✘
[bai2024curveformer++] 2024 The approach models the 3D detection task as a curve propagation problem, where each lane is represented by a curve query with a dynamic and ordered anchor point set. OpenLane, ONCE-3DLanes. N/A Transformer           ✘
@@
TABLE III: A brief introduction to the datasets in monocular 3D lane line detection, including their source papers, parent dataset, data type, resolution, data size, the year in which they were created, availability, and the link to their open-source repositories.
@@
OpenLane [chen2022persformer] Waymo [Sun_2020_CVPR] Real-world 1920 × 1280 200K 2022 link
@@
LLAMAS-3DLD [kim2023d] LLAMAS [9022318] Real-world 1276 × 717 6.6K 2023 -
@@ IV-D Other Datasets
The paper “3D-LaneNet: End-to-End 3D Multiple Lane Detection” [garnett20193d] presents two distinct datasets: the Synthetic-3D-Lanes dataset and the 3D-Lanes dataset. The Synthetic-3D-Lanes dataset, created via the open-source graphics engine, Blender, consists of 300K training and 5K testing examples, each featuring a 360×480 pixel image alongside associated ground truth parameters such as 3D lanes, camera height, and pitch. Exhibiting substantial diversity in lane topography, object placement, and scene rendering, this dataset provides a valuable resource for method development and ablation studies. Complementarily, the 3D-Lanes dataset is a real-world ground truth labeled data collection, compiled through the utilization of a multi-sensor setup - a forward-looking camera, a Velodine HDL32 lidar scanner, and a high-precision IMU. This dataset consists of six independent driving records, each recorded on a different road segment, totaling nearly two hours of driving time. With the aid of Lidar and IMU data, aggregated lidar top view images were generated and used alongside a semi-manual annotation tool to establish the ground truth. In total, 85,000 images were annotated, of which 1,000, derived from a separate drive, were designated as the test set, with the remainder serving as the training set. The 3D-Lanes dataset proves instrumental in validating the transferability of the proposed approach to real-world data and enabling qualitative analysis. While the Synthetic-3D-Lanes dataset is made available to the research community, the real-world 3D-Lanes dataset remains proprietary and is not publicly accessible. It’s worth noting that despite its availability, the Synthetic-3D-Lanes dataset has not seen wide adoption for benchmarking in subsequent studies in the field.
@@
The datasets introduced above encompass a variety of publicly available road scenes. Current mainstream research primarily focuses on favorable daytime scenes that are suitable for and conducive to three-dimensional lane detection, with sufficient lighting and favorable weather conditions. Many car companies and original equipment manufacturers in the industry possess a large volume of data, however, they are reluctant to share it publicly due to concerns related to intellectual property, industrial competition, and General Data Protection Regulations (GDPR). As a result, the lack of sufficient annotated data for accurately understanding dynamic weather conditions, such as nighttime, smoggy weather, and edge cases, remains a challenging task in autonomous driving research.
This research area is among the challenges that the community has not yet adequately addressed. In this section, we present critical viewpoints on the current state of three-dimensional lane detection in autonomous driving, summarize a series of challenges, and provide research direction recommendations to help the community make further progress and effectively overcome these challenges.

### v3

@@ II MONOCULAR 3D LANE DETECTION IN AD
Balanced Scene 3DLaneNet [garnett20193d] 91 RTX 3080Ti 89.3 86.4 0.068 0.477 0.015 0.202
Gen-LaneNet [guo2020gen] 102 RTX 3080Ti 90.1 88.1 0.061 0.496 0.012 0.214
CLGo [liu2022learning] 128 RTX 3080Ti 94.2 91.9 0.061 0.361 0.029 0.250
PersFormer [chen2022persformer] 29 RTX 3080Ti - 92.9 0.054 0.356 0.010 0.234
Anchor3DLane [huang2023anchor3dlane] 148 RTX 3080Ti 97.2 95.6 0.052 0.306 0.015 0.223
BEV-LaneDet [wang2023bev] 158 RTX 3080Ti - 98.7 0.016 0.242 0.020 0.216
LATR [luo2023latr] 14 RTX 3080Ti 97.9 96.8 0.022 0.253 0.007 0.202
D-3DLD [kim2023d] 22.6 Tesla A100 95.5 93.3 0.039 0.412 0.011 0.234
Rarely Observed 3DLaneNet [garnett20193d] 91 RTX 3080Ti 74.6 72.0 0.166 0.855 0.039 0.521
Gen-LaneNet [guo2020gen] 102 RTX 3080Ti 79.0 78.0 0.139 0.903 0.030 0.539
CLGo [liu2022learning] 128 RTX 3080Ti 88.3 86.1 0.147 0.735 0.071 0.609
PersFormer [chen2022persformer] 29 RTX 3080Ti - 87.5 0.107 0.782 0.024 0.602
Anchor3DLane [huang2023anchor3dlane] 148 RTX 3080Ti 96.9 94.4 0.094 0.693 0.027 0.579
BEV-LaneDet [wang2023bev] 158 RTX 3080Ti - 99.1 0.031 0.594 0.040 0.556
LATR [luo2023latr] 14 RTX 3080Ti 97.3 96.1 0.050 0.600 0.015 0.532
Visual Variations 3D-LaneNet [garnett20193d] 91 RTX 3080Ti 74.9 72.5 0.115 0.601 0.032 0.230
Gen-LaneNet [guo2020gen] 102 RTX 3080Ti 87.2 85.3 0.074 0.538 0.015 0.232
CLGo [liu2022learning] 128 RTX 3080Ti 89.2 87.3 0.084 0.464 0.045 0.312
PersFormer [chen2022persformer] 29 RTX 3080Ti - 89.6 0.074 0.430 0.015 0.266
Anchor3DLane [huang2023anchor3dlane] 148 RTX 3080Ti 93.6 91.4 0.068 0.367 0.020 0.232
BEV-LaneDet [wang2023bev] 158 RTX 3080Ti - 96.9 0.027 0.320 0.031 0.256
LATR [luo2023latr] 14 RTX 3080Ti 96.6 95.1 0.045 0.315 0.016 0.228
TABLE III: An in-depth comparison of various methods on the ApolloSim dataset under three different split settings. “N” and “F” are short for near and far, respectively.
3D-LaneNet [garnett20193d] 91 RTX 3080Ti 44.1 40.8 46.5 47.5 41.5 32.1 41.7
GenLaneNet [guo2020gen] 102 RTX 3080Ti 32.3 25.4 33.5 28.1 18.7 21.4 31.0
PersFormer [chen2022persformer] 29 RTX 3080Ti 50.5 42.4 55.6 48.6 46.6 40.0 50.7
Anchor3DLane [huang2023anchor3dlane] 148 RTX 3080Ti 54.3 47.2 58.0 52.7 48.7 45.8 51.7
BEV-LaneDet [wang2023bev] 158 RTX 3080Ti 58.4 48.7 63.1 53.4 53.4 50.3 53.7
GroupLane [li2023grouplane] 38.2 RTX 2080 64.1 - - - - - -
LATR [luo2023latr] 14 RTX 3080Ti 61.9 55.2 68.2 57.1 55.4 52.3 61.5
DecoupleLane [han2023decoupling] - - 51.2 43.5 57.3 - 48.9 43.5 -
TABLE IV: Comparison of various 3D methods on the OpenLane benchmark, evaluating the F-Score on both the entire validation set and separate sets corresponding to five different scenarios.
Method FPS Device F1(%) Precision(%) Recall(%) CD Error(m)
3D-LaneNet [garnett20193d] 91 RTX 3080Ti 44.73 61.46 35.16 0.127
Gen-LaneNet [guo2020gen] 102 RTX 3080Ti 45.59 63.95 35.42 0.121
SALAD [yan2022once] - - 64.07 75.90 55.42 0.098
PersFormer [chen2022persformer] 29 RTX 3080Ti 74.33 80.30 69.18 0.074
WS-3D-Lane [ai2022ws] 26.7 RTX 3080Ti 77.02 84.51 70.75 0.058
Anchor3DLane [huang2023anchor3dlane] 148 RTX 3080Ti 74.87 80.85 69.71 0.060
GroupLane [li2023grouplane] 38.2 RTX 2080 80.73 82.56 78.90 0.053
LATR [luo2023latr] 14 RTX 3080Ti 80.59 86.12 75.73 0.052
DecoupleLane [han2023decoupling] - - 75.07 81.19 69.26 0.062
TABLE V: The performance of 3D lane detection on the ONCE-3DLane dataset. CD error represents the Chamfer distance between the predicted lane line and the ground truth.
@@ III PERFORMANCE EVALUATION OF 3D LANE DETECTION
This section will discuss the performance evaluation of monocular 3D lane detection models. Herein, we explain the evaluation metrics, different types of objective functions, analyze the computational complexity, and finally provide quantitative comparisons of various models. The nomenclature of the used notations is given in Table II.
First, we present the visualization results of 3D lane line detection. As some algorithms are not open-sourced, we have only performed visualization tests on the ApolloSim dataset and OpenLane dataset using some open-source algorithms. These algorithms have been trained on the ApolloSim dataset and OpenLane dataset, and the visualization results are shown in Fig. 8 and Fig. 9, where the red lines represent the predicted lane lines, and the blue lines represent the ground truth lane lines. Next, we will introduce the evaluation metrics, the loss function used for training the algorithms, as well as the quantitative test results for 3D lane detection on public datasets, including performance metrics and a comparison of processing delays.
@@ III-A Evaluation Metrics for 3D Lane Detection
Building only predictive monocular 3D lane detection models is not a wise and trustworthy decision for safe AD unless it is tested on unseen data. Most models evaluate their performance on a disjoint set of the same dataset used for training, i.e., the test data remains new to the training model. Deep learning models for monocular 3D lane detection are evaluated using some general metrics to assess the best results based on true values. For monocular 3D lane detection tasks, different types of evaluation metrics are available for these tasks, which we review next as follows:
@@ III-A7 Z Error
III-A8 CD Error
$CD=\frac{1}{|P|}\sum\limits_{p\in P}\min\limits_{q\in Q}||p-q||^{2}+\frac{1}{|Q|}\sum\limits_{q\in Q}\min\limits_{p\in P}||q-p||^{2}$ (8)
@@ III-C Quantitative Analysis of Monocular 3D Lane Detection Models
This section elaborates on the quantitative empirical analysis of monocular 3D lane detection methods surveyed in this paper. For quantitative assessment, we utilize four evaluation metrics to examine the performance of each monocular 3D lane detection method on the ApolloSim dataset: AP, F-Score, x error and z error, and we report the results in Table III. On the OpenLane dataset, we evaluated the F-Score of each model, as shown in Table IV. On the ONCE-3DLane dataset, we evaluate four metrics, namely: F-Score, Precision, Recall, and CD error, the results are reported om Table V. The qualitative visualization results on the ApolloSim dataset and OpenLane dataset are shown in Fig. 8 and Fig. 9.
Additionally, computational efficiency is taken into account by reporting the Frames Per Second (FPS) achievable by each method during the inference process. The overall runtime of these models is reported in Table III, IV, and V. When evaluating computational efficiency, we use our unified computing platform to measure latency for open-source algorithms. For methods that are not open-source, we directly utilize the reported latency and corresponding computing platform details provided in the original papers. Our experimental platform’s CPU configuration includes an Intel(R) Core i9-12900K CPU processor running on the Ubuntu 20.04 operating system, while the GPU used in the experiments is an NVIDIA GeForce RTX 3080Ti GPU with 12 GB of graphics memory. From the table above, it can be observed that all methods have the potential to operate in real-time within AD systems, thus affirmatively addressing the question we raised in Section I-B. Among these methods, BEV-LaneDet [wang2023bev], Anchor3DLane [huang2023anchor3dlane], and CLGo [liu2022learning] demonstrate very fast processing speeds, which is advantageous for practical deployment. Methods like LATR [luo2023latr], D-3DLD [kim2023d], and WS-3D-Lane [ai2022ws], while meeting real-time requirements, exhibit higher processing latencies and would benefit from acceleration techniques, such as utilizing TensorRT for optimized deployment.
@@ IV DATASETS
In deep learning-based vision tasks, an equally important component is the dataset. In this section, we introduce the datasets used for monocular 3D lane detection tasks. Some of these datasets are open-sourced and popular, frequently used by the community, while others are only described in papers and are not open-sourced. Whether open-source or proprietary, we have compiled a detailed table to provide a more intuitive overview of these datasets, as presented in Table VI.
Figure 8: Visualization results of several open-source algorithms on the ApolloSim dataset [guo2020gen], with 3D-LaneNet [efrat20203d], Gen-LaneNet [guo2020gen], CLGo [liu2022learning], WS-3D-Lane [ai2022ws], and Anchor3DLane [huang2023anchor3dlane] algorithms displayed from top to bottom. In the images, the blue lines represent the ground truth of the lane lines, while the red lines represent the lane lines predicted by the algorithms.
Figure 9: Visualization results of several open-source algorithms on the OpenLane dataset [chen2022persformer], with Gen-LaneNet [guo2020gen], 3D-LaneNet [efrat20203d], Persformer [chen2022persformer], Anchor3DLane [huang2023anchor3dlane] and LATR [luo2023latr] algorithms displayed from top to bottom. In the images, the blue lines represent the ground truth of the lane lines, while the red lines represent the lane lines predicted by the algorithms.
@@ IV-A Apollo 3D Lane Synthetic Dataset
IV-B OpenLane
TABLE VI: A brief introduction to the datasets in monocular 3D lane line detection, including their source papers, parent dataset, data type, resolution, data size, the year in which they were created, availability, and the link to their open-source repositories.
@@ IV-B OpenLane
OpenLane [chen2022persformer] Waymo [sun2020scalability] Real-world 1920 × 1280 200K 2022 link
@@
LLAMAS-3DLD [kim2023d] LLAMAS [behrendt2019unsupervised] Real-world 1276 × 717 6.6K 2023 -
@@ IV-D Other Datasets
In [garnett20193d], it presents two distinct datasets: the Synthetic-3D-Lanes dataset and the 3D-Lanes dataset. The Synthetic-3D-Lanes dataset, created via the open-source graphics engine, Blender, consists of 300K training and 5K testing examples, each featuring a 360×480 pixel image alongside associated ground truth parameters such as 3D lanes, camera height, and pitch. Exhibiting substantial diversity in lane topography, object placement, and scene rendering, this dataset provides a valuable resource for method development and ablation studies. Complementarily, the 3D-Lanes dataset is a real-world ground truth labeled data collection, compiled through the utilization of a multi-sensor setup - a forward-looking camera, a Velodine HDL32 lidar scanner, and a high-precision IMU. This dataset consists of six independent driving records, each recorded on a different road segment, totaling nearly two hours of driving time. With the aid of Lidar and IMU data, aggregated lidar top view images were generated and used alongside a semi-manual annotation tool to establish the ground truth. In total, 85,000 images were annotated, of which 1,000, derived from a separate drive, were designated as the test set, with the remainder serving as the training set. The 3D-Lanes dataset proves instrumental in validating the transferability of the proposed approach to real-world data and enabling qualitative analysis. While the Synthetic-3D-Lanes dataset is made available to the research community, the real-world 3D-Lanes dataset remains proprietary and is not publicly accessible. It’s worth noting that despite its availability, the Synthetic-3D-Lanes dataset has not seen wide adoption for benchmarking in subsequent studies in the field.
As shown in Table VI, although there are a number of datasets currently available in the field of 3D lane detection, only a few are open-source. Regarding the question raised in section I-B, we believe that while these datasets do include some complex scenarios, they remain insufficient. For example, there is still a lack of datasets under nighttime and extreme weather conditions. To promote further development in this field, more challenging data should be added in the future. In addition to collecting real-world data, using the latest text-to-image and text-to-video generation technologies also offers promising possibilities for gathering data on challenging scenarios.
@@
The datasets introduced above encompass a variety of publicly available road scenes. Current mainstream research primarily focuses on favorable daytime scenes that are suitable for and conducive to 3D lane detection, with sufficient lighting and favorable weather conditions. Many car companies and original equipment manufacturers in the industry possess a large volume of data, however, they are reluctant to share it publicly due to concerns related to intellectual property, industrial competition, and General Data Protection Regulations (GDPR). As a result, the lack of sufficient annotated data for accurately understanding dynamic weather conditions, such as nighttime, smoggy weather, and edge cases, remains a challenging task in AD research.
This research area is among the challenges that the community has not yet adequately addressed. In this section, we present critical viewpoints on the current state of 3D lane detection in AD, summarize a series of challenges, and provide research direction recommendations to help the community make further progress and effectively overcome these challenges.